Let 𝑀𝑋 (𝑑) = π‘˜π‘’ 𝑑 1βˆ’0.8𝑒 𝑑 with Β΅=5 and define 𝑍 = 2𝑋 +3 , find Var(Z).


1
Expert's answer
2021-12-30T17:40:35-0500

Solution:

"M_X(t)=ke^{t\left(1-0.8e^t\right)}, \mu=5,Z=2X+3"

"\Rightarrow M'_X(t)=ke^{t\left(1-0.8e^t\right)}\left(1-0.8e^t-0.8e^tt\right)
\\E(X)=\mu=5\ (given)
\\E(X)=M'_X(0)"

So, "k(1)\left(1-0.8(1)-0.8(1)(0)\right)=5"

"\Rightarrow k(0.2)=5
\\ \Rightarrow k=25"

So, we have, "M_X(t)=25e^{t\left(1-0.8e^t\right)}"

Now, "M''_X(t)=25[0.64e^{3t-0.8e^tt}t^2-2.4e^{2t-0.8e^tt}t+1.28e^{3t-0.8e^tt}t-3.2e^{2t-0.8e^tt}+0.64e^{3t-0.8e^tt}+e^{t\left(-0.8e^t+1\right)}]"

"M''_X(0)=E(X^2)"

So, "M''_X(0)=25[0.64(0)-2.4(0)+1.28(0)+3.2(1)+0.64(1)+1]"

"=25[3.2+1.64]=121"

"Var(Z)=Var(2X+3)=4Var(X)+0=4Var(X)"

Also, "Var(X)=E(X^2)-[E(X)]^2=121-5^2=96"

Then, "Var(Z)=4(96)=384"